OpenStudy (anonymous):

f(x) = x(2x – 1) and g(x) = 2x – 1 f(x)/g(x)

OpenStudy (agreene):

\(f(x)=x(2x-1)\) \(g(x)=2x-1\) \[\frac{f(x)}{g(x)}=\frac{x(2x-1)}{2x-1}\] then you would simplify.

OpenStudy (anonymous):

what would it be after you simplify

OpenStudy (agreene):

what do you think it would be?

OpenStudy (anonymous):

1/x or x/1?

OpenStudy (agreene):

x/1 is correct, but that too can be simplified

OpenStudy (anonymous):

x

OpenStudy (anonymous):

thank you

OpenStudy (agreene):

no problem